# Exploring Concept Depth: How Large Language Models Acquire Knowledge at Different Layers?

### v1

@@ 4.2 Datasets
HateEval [21]: HateEval has English tweets which were annotated hierarchically.
STSA [17]: STSA includes movie reviews, half of which were considered positive and the other half negative. Each label is extracted from a longer movie review and reflects the writer’s overall intention for this review.
IMDb [20]: IMDb is a benchmark dataset for binary sentiment classification.
Sarcasm [25]: Sarcasm is a high-quality news headlines dataset that is annotated as sarcastic or not sarcastic.
StrategyQA [11]: StrategyQA contains questions across all knowledge domains to elicit creative and diverse yes/no questions that require implicit reasoning steps.
Coinflip [34]: Coinflip includes questions about coin flipping. This task requires the model to determine if a coin remains heads up after it is either flipped or left unflipped by individuals.
4.3 The Robustness of Internal Representations
4.3.1 Adding Noise
To quantify the robustness of the LLM with respect to its internal representation, when we input these questions into the LLM, we add a random string of noise in front of the input question to interfere the inference prosecc. For instance, we transform a given input question $q=\{q_{1},...,q_{i}\}$ into $q^{\prime}=\{N,q_{1},...,q_{i}\}$. Here $q_{i}$ represents the single sentence in question $q$ and $N$ is the noise string. The noise, which disrupts the classification tasks, originates from two separate random strings. These strings are inserted at the beginning of the sentence with an equal probability of 50%.
4.3.2 Quantization Settings
Quantization has been a rule of thumb for faster LLM inference. The general fact is that by using lower precision level allows LLMs to be run on less capable hardware with an acceptable reduction in their ability and accuracy. Our approach to applying quantization is to explore its effect on Concept Depth, whether it still maintains the previous Concept Depth or slows down the process of understanding knowledge. Quantization is set as 8-bit, 16-bit, and full 32-bit.
4.4 Metrics for Accuracy Variation
Definition 1 (Variation Rate) Given an LLM probing classifier $M=\{q,y,z,d\}$ ($q$, $y$, $z$, and $d$ are the input question, ground truth binary label, predicted label and total amount of layers, respectively), it has the accuracy at i-th layer $\alpha_{i}=\frac{1}{|z|}*\sum_{k=1}^{|z|}{[y_{k}=z_{k}]},i\in\{0,1,2,...,d-1\}$. We denote the variation rate $\beta_{i}=\alpha_{i}/\alpha_{i-1},i\in\{1,2,...,d-1\}$.
We introduce two metrics to capture variations in accuracy: the jump point and the converging point and define them by the given the definition of variation rate.
Definition 2 (Jump Point) We denote $J(M,D)=\min\{\frac{i}{d}\}$ s.t. $\beta_{i}>=1.1$, $i\in\{1,2,...,d-1\}$, as the jump point, where M and $D=(q,y)$ represents the LLM classifier and the dataset.
When a noticeable boost in accuracy is observed, the jump point signals the model’s recognition of critical features or patterns within the dataset.
Definition 3 (Converging Point) We denote $C(M,D)=\max\{\frac{i}{d}\}$ s.t. $|\beta_{i}-1|<0.03$, $i\in\{1,2,...,d-1\}$, as the converging point, where $M$ and $D=(q,y)$ represents the LLM classifier and the dataset.
As the accuracy plateaus or begins to decline, the convergence point indicates the model’s learning saturation or its peak learning capacity from the dataset.
Analyzing these metrics provides deeper insight into the model’s learning dynamics and its adaptability to various data types.
@@ 5 Experimental Analysis
RQ2: Do different size LLMs but the same series (e.g., Gemma series) have consistent Concept Depth? (Section 5.2)
@@ 5.1 Comparison Among the Datasets
In this section, we delve into an evaluative comparison of performance across a range of datasets, utilizing Figure 4 to detail the layer-wise accuracy of all nine LLMs over nine distinct datasets. Table 3 provides the accuracy, F1-score, and AUC metrics for parts of the layers (first, 25% depth, 50% depth, 67% depth, 83% depth, last) of each model over the nine datasets from the quantitative side. A performance threshold of 0.7 accuracy is applied to assess the models’ effective comprehension of concepts. This examination leads to two general observations. Firstly, regarding different concepts, LLMs exhibit varying accuracy trends across their layers; within individual concepts, however, different LLMs tend to display consistent accuracy patterns across these layers. Secondly, in tasks that require varying levels of conceptual understanding, the LLMs demonstrate their understanding across different neural network layers, indicating a layered approach to processing complex concepts.
Among the three factual concept datasets, significant variations in trends are observed across the models. The Cities dataset exhibits a sharp increase in comprehension at the lower layers, stabilizing in the higher layers, indicating a strong grasp of the concept. The CommonClaim dataset has become stable in early layers. On the other hand, the accuracy improvement of the nine LLMs trained on Counterfact was relatively difficult to achieve, utilizing deeper layers, and the accuracy was lower than that of many other datasets. Therefore, we can conclude that Counterfact is a complex task compared to the other datasets.
In datasets centered on emotional concept comprehension (STSA, IMDb, Sarcasm, and HateEval), despite varying levels of understanding, all models demonstrate a rise in accuracy at the initial layers, with convergence occurring in the intermediate layers. Although HateEval essentially reaches stability in the second or third layer, its accuracy reaches up to 0.8. This pattern suggests that LLMs primarily aggregate representations from lower layers to grasp emotional concepts.
Meanwhile, the datasets StrategyQA and Coinflip, which demand specific reasoning skills, tend to display a bell-shaped accuracy trajectory in all models, with peak accuracy observed in the middle layers. Such patterns underscore the intricate complexity associated with reasoning tasks.
The performance of classification tasks can be categorized into three types. 1) For Cities, STSA, IMDb, and Sarcasm, the LLMs suddenly understand the tasks at intermediate layers. 2) For CommonClaim and HateEval, the LLMs have already understood the tasks in shallower layers. 3) For Counterfact, StrategyQA, and Coinflip, The tasks are more difficult to understand, compared with others. Therefore, we consider the tasks in type 1 and 2 easy tasks, and those in type 3 are complex tasks.
@@ 5.2 Comparison Among the Number of Parameters
This section offers a comparative analysis of LLMs within their respective families, examining both accuracy levels and converging points across the models. Figure 3 reveals two recurring patterns within these families: for the tasks whose accuracy improved significantly by model learning, larger models tend to show converging points at earlier layers, suggesting they achieve their own peak comprehensions of concepts at lower layers; for the tasks whose accuracy do not change a lot, all LLM models show the converging points at early layers.
Two notable exceptions to this trend appear in the QWen family, over the Coinflip and IMDb datasets. For Coinflip, larger models exhibit delayed convergence. This deviation underscores the complexity of the reasoning required, illustrating how this task challenges even the larger models to extend their depth of understanding further. For IMDb, converging points fluctuate with the increasing size of the model because the number of layers is different among different sizes of LLMs, which amplifies the differences. These exceptions are also found in the Gemma family.
Furthermore, in Figure 3, we explore the peak accuracy levels across all layers for LLMs of differing sizes. The overarching trend indicates that larger models consistently achieve superior peak performance. This observation not only supports the notion that scaling up models enhances their effectiveness but also suggests that larger models develop more robust internal representations, validating the benefits of training models with greater capacity.
By comparing different sizes of models from the same LLM family, we have two observations. 1) As the number of parameters increases, peak accuracy gradually increases, and the converging point gradually advances. 2) Larger models grasp the concepts earlier and better.
@@ 5.3 Comparison Among the LLM Families
In this section, we examine how LLMs from various families, possessing a similar parameter count, process concepts as reflected by their converging points and peak accuracies. The overarching trends are highlighted in Figure 6, with detailed statistics on a layer-by-layer basis provided in Figure 5. Our findings reveal that while LLMs across different families may reach nearly identical peak accuracies, the layers at which they converge to these peaks can vary. For instance, in the HateEval and Counterfact datasets, we observe models converging at significantly deeper layers. This variation suggests that despite similar parameter scales, different models may employ varied mechanisms to tackle the same problems, reflecting the diversity in how models interpret and process complex information.
With the same magnitude of model parameters, the datasets generally perform comparably. The exceptions are Coinflip and STSA.
Figure 7: Noise and precision evaluation on STSA and Sarcasm of Gemma-2b.
@@ 5.4 Ablation Study
In this section, we investigate the impact of noise and precision reduction on the learning classification task of LLMs. For noise perturbation, we randomly add two short strings to the beginning of each problem and observe how the LLM’s performance changes. We also evaluate language models for 8, 16, and full 32 bits for precision. For example, Figure 7 illustrates the effect of adding noise and reducing bit representations to the Gemma-2b model. The addition of noise causes the learning curve to shift to the right, indicating a reduction in the speed of convergence. This suggests that the presence of noise in the input data can negatively impact the LLM’s learning process, slowing down its ability to converge to an optimal solution. We also see that there isn’t much difference between 32 and 16 bits, and the convergence rate becomes slower when we reduce the model to 8 bits. Therefore, we can compress the LLMs to 16 bits for future designs.
6 Discussion
Probing technology has the potential to revolutionize various aspects of machine learning and accelerate inference in large models. One exciting application involves using probes to determine the layer at which a concept is learned within a neural network in a large language model. This information can then be leveraged to perform targeted model pruning, specifically removing the layers that follow the identified layer. This pruning process can significantly speed up the inference process in large language models.
By deploying probes throughout a neural network, we can gather insights into how information is processed and transformed as it flows through the network. These probes act as detectors, identifying the layers where specific concepts or features are learned. This knowledge allows us to identify the critical layers that contribute significantly to the model’s final predictions.
Once the layer responsible for learning a particular concept is identified, targeted model pruning can be applied. Pruning involves removing unnecessary or less influential layers from the network while preserving its overall performance. By cutting off the layers that follow the identified layer, we eliminate computational overhead and reduce the model’s complexity, resulting in faster inference times.
@@ 7 Conclusions
This paper proposes the Concept Depth, the phenomenon that different concepts are learned in different layers of large language models, i.e. more difficult concepts are fully acquired with deeper layers. We conducted a number of experiments around Concept Depth using probe techniques. Our research suggests that LLMs tend to effectively categorize easy tasks, indicating that these concepts are learned in the first few initial layers. In contrast, complex tasks may only be recognizable (if at all) in deeper layers, and LLMs of the same size perform largely consistently across datasets in terms of concept depth.
@@ A.1 Metrics for Parts of the Layers
Gemma-2b (18 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
Gemma-7b (28 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
LlaMA-7b (32 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
LlaMA-13B (40 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
QWen-0.5B (24 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
QWen-1.8B (24 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
QWen-7B (40 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
QWen-14B (32 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
QWen-14B (40 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
A.2 Examples of Each Dataset
@@
A.3 LLM structure

### v4

@@ 5 Experimental Analysis
RQ2: Do different size LLMs in the same family (e.g., the LLaMA family) have consistent Concept Depth? (Section 5.2)
@@ 5.1 Comparison Among the Datasets
We delve into an evaluative performance comparison across a range of datasets, utilizing Figure 3 to detail the layer-wise accuracy of all nine LLMs over nine distinct datasets. Table 3 shows the detailed numerical results for Figure 3, as well as the F1-score and AUC. A performance threshold of 0.7 accuracy is applied to assess the models’ effective comprehension of concepts. This examination leads to two general observations. Firstly, regarding different concepts, LLMs exhibit varying accuracy trends across their layers. For example, Cities approaches perfect accuracy fast; in contrast, datasets requiring high-level reasoning such as StrategyQA will not reliably converge to accuracy above 0.7, indicating that they have different “Concept Depth”. Within individual concepts, however, different LLMs tend to display consistent accuracy patterns across these layers. Secondly, in tasks that require varying levels of conceptual understanding, the LLMs demonstrate their understanding across different layers, indicating a layered approach to processing complex concepts.
Significant variations in trends are observed across the models among the three factual concept datasets. Cities exhibits a sharp increase in comprehension at lower layers, stabilizing in higher layers, indicating a strong grasp of the concept. CommonClaim has become stable in early layers. Besides, the accuracy improvement of the nine LLMs trained on Counterfact was relatively difficult to achieve, utilizing deeper layers, and the accuracy was lower than that of many other datasets. Therefore, we can conclude that Counterfact is more complex.
In datasets centered on emotional concept comprehension (STSA, IMDb, Sarcasm, and HateEval), despite varying levels of understanding, all models demonstrate a rise in accuracy at the initial layers, with convergence occurring in the intermediate layers. Although HateEval essentially reaches stable at the initial layers, its accuracy reaches up to 0.8, suggesting that LLMs primarily aggregate representations from lower layers to grasp emotional concepts. Meanwhile, StrategyQA and Coinflip, which demand specific reasoning skills, tend to display a bell-shaped accuracy trajectory in all models, with peak accuracy observed in the middle layers. Such patterns underscore the intricate complexity associated with reasoning tasks.
Remark 1 We categorize the performances into three types. 1) For Cities, STSA, IMDb, and Sarcasm, the LLMs suddenly understand the tasks at intermediate layers. 2) For CommonClaim and HateEval, the LLMs have already understood the tasks in shallower layers. 3) For Counterfact, StrategyQA, and Coinflip, The tasks are more difficult to understand compared with others. Therefore, we consider the tasks in type 1 and 2 easy tasks, and those in type 3 are complex.
@@ 5.2 Comparison Among the Number of Parameters
This section offers a comparative analysis of LLMs within their respective families, examining both accuracy levels and converging points across the models. Figure 4 reveals two recurring patterns within these families: for tasks with accuracy improves dramatically by model learning, larger models tend to show converging points at earlier layers, suggesting they achieve their own peak comprehensions of concepts at lower layers; for tasks with accuracy changes little, all LLMs show the converging points at early layers.
Two notable exceptions to this trend appear in the Qwen family over the Coinflip and IMDb datasets. For Coinflip, larger models exhibit delayed convergence. This deviation underscores the complexity of the reasoning required, illustrating how this task challenges even the larger models to extend their depth of understanding further. For IMDb, converging points fluctuate with the increasing size of the model because the number of layers is different among different sizes of LLMs, which amplifies the differences. These exceptions are also found in the Gemma family.
Furthermore, in Figure 4, we explore the peak accuracy levels across all layers for LLMs of differing sizes. The overarching trend indicates that larger models consistently achieve superior peak performance. This observation not only supports that scaling up models enhances their effectiveness but also suggests that larger models develop more robust internal representations, validating the benefits of training models with greater capacity.
Remark 2 We have two observations by comparing different sizes of models from the same LLM family. 1) As the number of parameters increases, peak accuracy gradually increases, and the converging point gradually advances. 2) Larger models grasp the concepts earlier and better.
@@ 5.3 Comparison Among the LLM Families
We examine how LLMs from various families, possessing a similar parameter count, process concepts as reflected by their converging points and peak accuracies. The overarching trends are highlighted in Figure 6, with detailed statistics on a layer-by-layer basis provided in Figure 5. Our findings reveal that while LLMs across different families may reach nearly identical peak accuracies, the layers at which they converge to these peaks can vary. For instance, in the HateEval and Counterfact datasets, we observe models converging at significantly deeper layers. This variation suggests that despite similar parameter scales, different models may employ varied mechanisms to tackle the same problems, reflecting the diversity in how models interpret and process complex information.
Remark 3 With the same number of model parameters, the models generally have a comparable understanding of the datasets.
@@ 5.4 Ablation Study
To quantify the robustness of the LLMs concerning their internal representation, we conducted ablation studies on noise perturbation and bit quantization. The result shows that adding noises or reducing model weights to 8 bits can make the accuracy converge slower. Compressing the LLMs to 16 bits doesn’t harm the understanding process too much. Details can be found in Section A.3.
@@ 6 Conclusions
This paper proposes Concept Depth, the phenomenon that different concepts are learned in different layers of LLMs, i.e., more difficult concepts are fully acquired with deeper layers. We conducted several experiments around Concept Depth using probe techniques. Our research suggests that LLMs tend to effectively categorize easy tasks, indicating that these concepts are learned in the first few initial layers. In contrast, complex tasks may only be recognizable (if at all) in deeper layers, and LLMs of the same size perform largely consistently across datasets regarding Concept Depth. Compressing the model weight to 16-bit representations for future LLMs’ designs is also a promising method for saving computation memory.
7 Limitations
The paper presents several opportunities for further exploration. Firstly, the datasets employed might not encompass the full spectrum of language tasks, offering a chance to expand the scope of the findings in a multilingual environment. Secondly, We did not experiment with very large open-source language models, thus allowing future researchers to investigate how scaling up the model size affects concept acquisition across different layers and enhances robustness. Moreover, we should also try different kinds of classifiers, including nonlinear models and neural network-based classifiers, to acquire more profound insights into how LLM representations differ across layers. These aspects highlight promising directions for continued advancement in the field. We will continue to explore intermediate representations to help us better understand the inner side of LLMs, as this challenge may also be open to other researchers in this field.
@@ A.1 Metrics for Parts of the Layers
Gemma-2B (18 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
Gemma-7B (28 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
LlaMA-7B (32 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
LlaMA-13B (40 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
Qwen-0.5B (24 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
Qwen-1.8B (24 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
Qwen-7B (40 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
Qwen-14B (32 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
Qwen-14B (40 Layers) StrategyQA Coinflip Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
@@
A.2 Description of the dataset
Cities Marks and Tegmark (2023): consists of statements about the location of cities and their veracity labels (e.g., The city of Zagreb is in Japan, which is wrong). We use 1496 of these samples.
CommonClaim Casper et al. (2023): A dataset of boolean statements, each labeled by two humans as common-knowledge-true, common-knowledge-false, or neither. We use 6000 of these samples.
Counterfact Meng et al. (2022): Counterfact includes myriad counterfactuals that allows quantitative testing of specificity and generalization when learning a counterfactual. We use 4000 of these samples.
HateEval Manolescu et al. (2019): HateEval has English tweets which were annotated hierarchically. We use 6000 of these samples.
STSA Kim (2014): STSA includes movie reviews, half of which were considered positive and the other half negative. Each label is extracted from a longer movie review and reflects the writer’s overall intention for this review. We use 6920 of these samples.
IMDb Maas et al. (2011): IMDb is a benchmark dataset for binary sentiment classification. We use 2000 of these samples.
Sarcasm Misra and Arora (2023): Sarcasm is a superior news headline dataset that tells if the headlines are sarcastic. We use 6000 of these samples.
StrategyQA Geva et al. (2021): StrategyQA contains questions across all knowledge domains to elicit creative and diverse yes/no questions that require implicit reasoning steps. We use 2290 of these samples.
Coinflip Wei et al. (2022b): Coinflip includes coin flipping queries, asking if a coin remains heads up after it is either flipped or left unflipped by individuals. We use 500 of these samples.
A.3 Ablation Study
Adding Noise. To quantify the robustness of the LLMs concerning their internal representation, when we input these questions into the LLM, we add a random string of noise in front of the question to interfere. For instance, we perturbate a question $q$ into $q^{\prime}=N\oplus q$. Here, $\oplus$ is the concatenate operation, and the noise string $N\in\{S_{1},S_{2}\}$ disrupts the classification tasks, satisfying
$P(N=S_{1})=P(N=S_{2})=50\%$
The probability distribution of $N$ is unrelated to their labels. The following is an example of STSA with a positive label.
Before adding noise:
After adding noise:
aaa The production values are of the highest and the performances attractive without being memorable. The sentence above is a movie review and reflects the writer’s overall intention for this review. According to the sentence, judge whether the emotion is Positive or Negative.
aaa can be substituted by bbb, both cases have equal probability.
Quantization Settings. Quantization has been a rule of thumb for faster LLMs’ inference. The general fact is that using a lower precision level allows LLMs to be run on less capable hardware with an acceptable reduction in their ability and accuracy. Our approach to applying quantization is to explore its effect on Concept Depth, whether it still maintains the previous Concept Depth or slows down the process of understanding knowledge. We quantize the model weights to 8, 16, and 32 bits.
Results. Figure 7 illustrates the effect of adding noise and reducing bit representations to the Gemma-2B model. The addition of noise causes the learning curve to shift to the right, indicating a reduction in the converging speed. This suggests that the presence of noise in the input data can hinder the LLM’s learning, slowing down its ability to converge to an optimal solution. We also see that there isn’t much difference between 32 and 16 bits, and the convergence rate slows when we reduce the model to 8 bits. Therefore, we may compress the LLMs to 16 bits for future designs.
Remark 4 Noises or 8-bit-quantization can cause the accuracy to converge more slowly. Compressing the LLMs to 16 bits doesn’t harm the understanding process too much. The layer-wise representations of LLMs are susceptible to noise and high-ratio quantization. Therefore, it is crucial to proceed cautiously when conducting high-ratio quantization inference.
Figure 7: Noise and quantization evaluation on STSA and Sarcasm of Gemma-2B.
A.4 Examples of Each Dataset
@@
Table 4: The category that each dataset belongs to.
Category Dataset
Fact Cities
Common
Counterfact
Emotion HateEval
STSA
IMDb
Sarcasm
Reasoning StrategyQA
Coinflip
A.5 LLM structure